OpenStudy (anonymous):

How many moles of sodium chloride (NaCl) solute are in 155 grams of an 85.5 percent by mass solution?

OpenStudy (anonymous):

my first reaction: (155 grams)* (0.855) / (formula weight of NaCl)

OpenStudy (anonymous):

i think that is wrong .

OpenStudy (anonymous):

what's the right answer? (how did you know it's wrong?)

OpenStudy (anonymous):

assuming i am reading the right textbook, percent by mass = (mass of solute ) / (mass of solution) so mass of NaCl = percent by mass * mass of solution = 0.855 * 155 grams so mol of NaCl = 0.855 * 155 grams / (FW of NaCl) and FW of NaCl is 22.99 + 35.45 grams / mol

OpenStudy (australopithecus):

looks good to me, I recommend doing the math one step at a time though to avoid error. (0.855 * 155 grams)/(22.99g/mol + 35.45 grams/mol) = moles of NaCl solute